40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 12 03:49:45 2011, Jenny, Summary, PSL, NPRO PDH-Locked to Ref Cav LIGO_block_diagram_2.pngopenlooperror.jpgIMG_3034.JPGclosedlooperror.jpg
    Reply  Fri Aug 12 20:33:57 2011, Dmass, Summary, PSL, NPRO PDH-Locked to Ref Cav TMP_INT_TF.pdf
Message ID: 5217     Entry time: Fri Aug 12 20:33:57 2011     In reply to: 5202
Author: Dmass 
Type: Summary 
Category: PSL 
Subject: NPRO PDH-Locked to Ref Cav 

To aid Jenny's valiant attempt to finish her SURF project, I did some things with the front end system over the last couple days, largely tricking Jamie into doing things for me lest I ruin the 40m RCG system. Several tribulations have been omitted.

We stole a channel in the frontend, in the proccess:

  1. Modified the C1GFD simulink model (now analog) to be "ADC -> TMP -> DAC" where TMP is a filter bank
    • C1GFD_TMP.adl (in /opt/rtcds/caltech/c1/medm/c1gfd) is the relevant part which connects the ADC to the DAC in the frontend
  2. Confirmed that the ADC was working by putting a signal in and seeing it in the frontend
  3. Could not get a signal out of the anti aliasing board
  4. Looked sad until Kiwamu found a breakout board for the SCSI cable coming from the DAC
  5. Used SR560 to buffer DAC output
    • drove a triangle wave with AWG into the TMP EXC channel (100 counts 1 Hz) and looked at it after the ~25 ft of BNC cable running between the DAC and the NRPO driver
    • wave looked funny (not like a triangle wave), maybe the DAC is not meant to push a signal so far, so added buffer
  6. Took the control signal going to the fast input of the NPRO driver (using the 500 Ohm SR560 output - see Jenny's diagram) and put it into the anti aliasing board of the ADC
  7. Added switchable integrator to filter bank with Foton
    • I couldn't get the names to display in the filter bank, so I looked sad again
    • Jamie and Koji both poked at the "no name displayed" problem but had no conclusions, so I decided to ignore it
    • I confirm that when the two filters were toggled "on" that the transfer function was as expected: simple integrator with a unity gain at ~10mHz - agrees with what Foton's Bode Plot tool says it should be (see attached DTT plot)
  8. I got Jamie to manually add the two epics channels from the TMP model to the appropriate .ini file so they would be recorded
    • C1:GFD-TMP_OUTPUT  (16 Hz)
    • C1:GFD-TMP_INMON    (16 Hz)
  9. RefCav heater servo seems to still be set up, so we can use existing channels:
    • C1:PSL-FSS_RCPID_SETPOINT (temp setpoint - will do +/-1C steps about 35 C)
    • C1:PSL-FSS_MINCOMEAS (In loop temp sensor - in C)
    • C1:PSL-FSS_RCTEMP (out of loop temp sensor - in C)
    • C1:PSL-FSS_TIDALSET (Voltage to heater - rails @ +/- 2V)
  10.  Closed loop on the control signal for the NPRO driver with an integrator, saw error signal go to zero
    • Turned up gain a little bit, saw some oscillations, then turned gain down to stop them, final gain = 2
  11. Left system on for Jenny to come in and do step responses
Attachment 1: TMP_INT_TF.pdf  8 kB  Uploaded Fri Aug 12 22:25:05 2011  | Hide | Hide all
TMP_INT_TF.pdf
ELOG V3.1.3-